# Pre-trained Model Guided Fine-Tuning for Zero-Shot Adversarial Robustness

### 29 Feb 2024 v2

@@ Abstract
Large-scale pre-trained vision-language models like CLIP have demonstrated impressive performance across various tasks, and exhibit remarkable zero-shot generalization capability, while they are also vulnerable to imperceptible adversarial examples. Existing works typically employ adversarial training (fine-tuning) as a defense method against adversarial examples. However, direct application to the CLIP model may result in overfitting, compromising the model’s capacity for generalization. In this paper, we propose Pre-trained Model Guided Adversarial Fine-Tuning (PMG-AFT) method, which leverages supervision from the original pre-trained model by carefully designing an auxiliary branch, to enhance the model’s zero-shot adversarial robustness. Specifically, PMG-AFT minimizes the distance between the features of adversarial examples in the target model and those in the pre-trained model, aiming to preserve the generalization features already captured by the pre-trained model. Extensive Experiments on 15 zero-shot datasets demonstrate that PMG-AFT significantly outperforms the state-of-the-art method, improving the top-1 robust accuracy by an average of 4.99%. Furthermore, our approach consistently improves clean accuracy by an average of 8.72%.
@@ 1 Introduction
Figure 1: Zero-shot robust accuracy (a) and clean accuracy (b) of CLIP and CLIPs fine-tuned on TinyImageNet [12] using various methods across multiple datasets. CLIP: Original CLIP. FT-Standard: CLIP fine-tuned on clean samples. FT-TeCoA: CLIP fine-tuned using [32]. PMG-AFT (ours): CLIP fine-tuned by our method.
@@ 2 Related Work
Finetuning and Catastrophic Overfitting. Fine-tuning is a prevalent strategy aimed at adapting pre-trained models to specific downstream tasks [13, 16]. However, when fine-tuning pre-trained vision-language models, one may encounter overfitting issues. During this process, the target model might deviate significantly from the pre-trained one, leading to overfitting on small-scale fine-tuning datasets [14, 49]. Within the adversarial training framework, the issue of overfitting still exists and is referred to as ”catastrophic overfitting” [39, 43]. In this situation, the model is influenced by the distribution of adversarial examples, leading to a significant decline in both robust accuracy and clean accuracy. Methods to address overfitting include linear interpolation [44], and parameter regularization [25]. These methods focus on parameter space to constrain the distance between two models while our method addresses overfitting by encouraging memory in the feature space.
@@ 3.3 PMG-AFT
Generalization Information Branch. As adversarial examples are generated based on a specific dataset, the robustness features acquired by the target model are confined and may excessively specialize to a particular downstream task dataset, consequently compromising generalizability. Our method introduces a generalization information branch to improve the generalization of adversarial robustness. The main body of this branch is an image encoder of the original pre-trained CLIP, which is denoted as $F_{ori}(\cdot)$. $I_{ori}$ denotes the embeddings of adversarial images it produces. Since the pre-trained model is a fixed deterministic function, the supervision objective of this branch essentially encourages the target model $F_{\theta}(\cdot)$ to output features that can predict the information of the original foundational model as much as possible, thereby helping to mitigate the problem of overfitting. We feed the adversarial examples $X_{a}$ into both the target model and the original pre-trained model, obtaining both prediction results based on the text embeddings $T$.
$P_{adv}$ $=\text{softmax}(I_{a}\cdot T^{\top}),$ (7)
$P_{ori}$ $=\text{softmax}(I_{ori}\cdot T^{\top}).$
@@ 4.1 Experimental Setup
Table 1: Adversarial zero-shot robust accuracies under PGD-10 [30] attack. We fine-tune the model on TinyImageNet [12] and evaluate six methods (rows) on 16 datasets (columns), presenting the accuracy for each dataset as well as the average accuracy, with the best results shown in bold. Here, CLIP represents the pre-trained CLIP model, while FT-standard refers to the model fine-tuned on clean datasets.
Method CIFAR10 CIFAR100 STL10 SUN397 Food101 Oxfordpet Flowers102 DTD EuroSAT Fgvc_Aircraft TinyImageNet ImageNet Caltech101 Caltech256 StanfordCars PCAM Average Time CLIP 42.18 17.57 67.77 10.49 3.28 8.98 4.88 18.75 3.84 0.39 0.39 6.09 15.82 23.76 2.34 51.61 17.38 0 FT-Standard 24.21 12.30 47.65 5.71 2.65 1.56 7.61 17.57 0.13 0.00 0.39 6.87 20.31 23.82 1.75 48.15 13.79 234 FT-TeCoA [32] 40.82 24.41 70.70 19.21 14.45 28.13 23.05 28.13 12.57 3.13 19.33 16.48 24.02 40.56 12.69 53.68 26.96 551 PMG-AFT (ours) 66.99 38.28 76.17 24.23 14.92 33.59 23.43 34.38 24.15 3.91 14.84 17.26 24.02 45.05 14.64 57.64 31.95 849 VP-TeCoA [32] 36.71 16.21 57.61 6.05 4.45 13.67 13.02 10.15 9.73 0.00 18.55 5.56 18.16 25.32 2.14 54.92 18.27 364 VPT-PMG-AFT (ours) 50.19 22.07 64.45 12.79 8.35 18.55 16.85 3.90 11.08 2.73 20.70 9.92 19.25 38.99 5.46 58.71 22.74 661
@@
Method CIFAR10 CIFAR100 STL10 SUN397 Food101 Oxfordpet Flowers102 DTD EuroSAT Fgvc_Aircraft TinyImageNet ImageNet Caltech101 Caltech256 StanfordCars PCAM Average Time CLIP 89.52 62.50 97.46 55.25 75.00 85.93 48.82 41.79 50.84 8.70 56.03 61.99 32.42 84.76 55.27 53.40 59.98 0 FT-Standard 88.67 58.78 95.89 46.61 70.07 75.00 41.60 41.79 38.60 8.57 58.59 51.28 29.10 79.88 37.10 54.74 54.76 234 FT-TeCoA [32] 66.79 41.01 89.25 47.01 52.81 70.31 36.13 35.94 18.88 7.81 48.83 43.67 28.32 72.98 37.89 37.89 46.99 551 PMG-AFT (ours) 83.98 58.39 92.97 56.41 66.40 84.76 42.96 41.02 35.28 6.25 46.87 56.75 30.46 82.94 48.24 48.24 55.71 849 VPT-TeCoA [32] 53.51 26.36 74.41 17.22 15.85 33.20 22.39 12.50 13.77 0.39 44.92 15.82 21.26 42.38 3.32 58.71 28.50 364 VPT-PMG-AFT (ours) 67.38 33.78 81.64 29.19 28.51 49.60 17.18 13.92 13.92 7.81 52.14 26.83 21.31 61.19 16.99 58.95 36.03 661
@@
Method CIFAR10 CIFAR100 STL10 SUN397 Food101 Oxfordpet Flowers102 DTD EuroSAT Fgvc_Aircraft TinyImageNet ImageNet Caltech101 Caltech256 StanfordCars PCAM Average Time CLIP 8.98 3.90 13.47 0.07 0.39 0.00 0.00 0.00 0.00 0.39 0.00 0.00 0.41 0.15 0.09 0.16 1.74 0 FT-TeCoA [32] 9.96 4.68 23.24 0.21 2.73 5.44 5.17 14.85 9.73 0.39 16.99 8.85 16.21 27.08 4.49 15.26 10.33 551 PMG-AFT (ours) 31.05 13.28 44.33 14.47 10.54 12.78 8.75 23.56 15.13 0.78 13.39 9.07 16.40 43.16 12.10 17.84 17.91 849
Implementation Details. We utilize the ViT-B/32 architecture of the CLIP model as the backbone and use the SGD optimizer to adversarially finetune the target model for 10 epochs. For fine-tuning, we update all parameters of the image encoder with a learning rate of 5e-5. For visual prompt, we introduce learnable parameters as prompts, adding them to both the image layer with the same dimensions as the original image and the token layer with a dimension of 100. The learning rate for the visual prompt is set at 40. We use $l_{\infty}$ norm PGD-2 [30] and PGD-10 [30] attacks with perturbation bounds of $\varepsilon=1/255,2/255$ and $4/255$ for both adversarial training and evaluation, respectively. For hyper-parameters, we set $\alpha$ to 1 and $\beta$ to 1. Each model included in this paper is trained on two NVIDIA GeForce RTX 3090 GPUs.
@@ 4.3 Performance against AutoAttack
Figure 4: Zero-shot adversarial robustness under different perturbation bounds ($\varepsilon=1/255,2/255$ and $4/255$). We employ the same perturbation bound for both training and testing. We present the robust accuracy (a) and the clean accuracy (b).
@@ 4.6 Ablation Study
Method Clean Avg Acc (%) Robust Avg Acc (%) Output + KL 55.71 31.95 Output + $L_{2}$ 48.21 27.23 Feature + $L_{2}$ 38.32 22.90 Feature + COS 52.98 24.53

### 10 Apr 2024 v3

@@ Abstract
Large-scale pre-trained vision-language models like CLIP have demonstrated impressive performance across various tasks, and exhibit remarkable zero-shot generalization capability, while they are also vulnerable to imperceptible adversarial examples. Existing works typically employ adversarial training (fine-tuning) as a defense method against adversarial examples. However, direct application to the CLIP model may result in overfitting, compromising the model’s capacity for generalization. In this paper, we propose Pre-trained Model Guided Adversarial Fine-Tuning (PMG-AFT) method, which leverages supervision from the original pre-trained model by carefully designing an auxiliary branch, to enhance the model’s zero-shot adversarial robustness. Specifically, PMG-AFT minimizes the distance between the features of adversarial examples in the target model and those in the pre-trained model, aiming to preserve the generalization features already captured by the pre-trained model. Extensive Experiments on 15 zero-shot datasets demonstrate that PMG-AFT significantly outperforms the state-of-the-art method, improving the top-1 robust accuracy by an average of 4.99%. Furthermore, our approach consistently improves clean accuracy by an average of 8.72%. Our code is available at here. [^1] https://github.com/serendipity1122/Pre-trained-Model-Guided-Fine-Tuning-for-Zero-Shot-Adversarial-Robustness
@@ 1 Introduction
Figure 1: Zero-shot robust accuracy (a) and clean accuracy (b) of CLIP and CLIPs fine-tuned on TinyImageNet [12] using various methods across multiple datasets. FT-Standard: CLIP fine-tuned on clean samples. FT-TeCoA: CLIP fine-tuned using [32]. PMG-AFT (ours): CLIP fine-tuned by our method.
@@ 2 Related Work
Finetuning and Catastrophic Overfitting. Fine-tuning is a prevalent strategy aimed at adapting pre-trained models to specific downstream tasks [13, 16]. However, when fine-tuning pre-trained vision-language models, one may encounter overfitting issues. During this process, the target model might deviate significantly from the pre-trained one, leading to overfitting on small-scale fine-tuning datasets [14, 49]. Within the adversarial training framework, the issue of overfitting still exists and is referred to as ”adversarial overfitting” [39, 43]. In this situation, the model is influenced by the distribution of adversarial examples, leading to a significant decline in both robust accuracy and clean accuracy. Methods to address overfitting include linear interpolation [44], and parameter regularization [25]. These methods focus on parameter space to constrain the distance between two models while our method addresses overfitting by encouraging memory in the feature space.
@@ 3.3 PMG-AFT
Generalization Information Branch. As adversarial examples are generated based on a specific dataset, the robustness features acquired by the target model are confined and may excessively specialize to a particular downstream task dataset, consequently compromising generalizability. Our method introduces a generalization information branch to improve the generalization of adversarial robustness. The main body of this branch is an image encoder of the original pre-trained CLIP, which is denoted as $F_{ori}(\cdot)$. $I_{ori-adv}$ denotes the embeddings of adversarial images it produces. Since the pre-trained model is a fixed deterministic function, the supervision objective of this branch essentially encourages the target model $F_{\theta}(\cdot)$ to output features that can predict the information of the original foundational model as much as possible, thereby helping to mitigate the problem of overfitting. We feed the adversarial examples $X_{a}$ into both the target model and the original pre-trained model, obtaining both prediction results based on the text embeddings $T$.
$P_{adv}$ $=\text{softmax}(I_{adv}\cdot T^{\top}),$ (7)
$P_{ori-adv}$ $=\text{softmax}(I_{ori-adv}\cdot T^{\top}).$
@@ 4.1 Experimental Setup
Table 1: Adversarial zero-shot robust accuracies under PGD-10 [30] attack. We fine-tune the model on TinyImageNet [12] and evaluate six methods (rows) on 16 datasets (columns), presenting the accuracy for each dataset as well as the average accuracy, with the best results shown in bold. The accuracy here is percentage.
Method CIFAR10 CIFAR100 STL10 SUN397 Food101 Oxfordpet Flowers102 DTD EuroSAT Fgvc_Aircraft TinyImageNet ImageNet Caltech101 Caltech256 StanfordCars PCAM Average Time (s) CLIP 42.18 17.57 67.77 10.49 3.28 8.98 4.88 18.75 3.84 0.39 0.39 6.09 15.82 23.76 2.34 51.61 17.38 0 FT-Standard 24.21 12.30 47.65 5.71 2.65 1.56 7.61 17.57 0.13 0.00 0.39 6.87 20.31 23.82 1.75 48.15 13.79 234 FT-TeCoA [32] 40.82 24.41 70.70 19.21 14.45 28.13 23.05 28.13 12.57 3.13 19.33 16.48 24.02 40.56 12.69 53.68 26.96 551 PMG-AFT (ours) 66.99 38.28 76.17 24.23 14.92 33.59 23.43 34.38 24.15 3.91 14.84 17.26 24.02 45.05 14.64 57.64 31.95 849 VP-TeCoA [32] 36.71 16.21 57.61 6.05 4.45 13.67 13.02 10.15 9.73 0.00 18.55 5.56 18.16 25.32 2.14 54.92 18.27 364 VPT-PMG-AFT (ours) 50.19 22.07 64.45 12.79 8.35 18.55 16.85 3.45 11.08 2.73 20.70 9.92 19.25 38.99 5.46 58.71 22.74 661
@@
Method CIFAR10 CIFAR100 STL10 SUN397 Food101 Oxfordpet Flowers102 DTD EuroSAT Fgvc_Aircraft TinyImageNet ImageNet Caltech101 Caltech256 StanfordCars PCAM Average Time (s) CLIP 89.52 62.50 97.46 55.25 75.00 85.93 48.82 41.79 50.84 8.70 56.03 61.99 32.42 84.76 55.27 53.40 59.98 0 FT-Standard 88.67 58.78 95.89 46.61 70.07 75.00 41.60 41.79 38.60 8.57 58.59 51.28 29.10 79.88 37.10 54.74 54.76 234 FT-TeCoA [32] 66.79 41.01 89.25 47.01 52.81 70.31 36.13 35.94 18.88 7.81 48.83 43.67 28.32 72.98 37.89 37.89 46.99 551 PMG-AFT (ours) 83.98 58.39 92.97 56.41 66.40 84.76 42.96 41.02 35.28 6.25 46.87 56.75 30.46 82.94 48.24 48.24 55.71 849 VPT-TeCoA [32] 53.51 26.36 74.41 17.22 15.85 33.20 22.39 12.50 13.77 0.39 44.92 15.82 21.26 42.38 3.32 58.71 28.50 364 VPT-PMG-AFT (ours) 67.38 33.78 81.64 29.19 28.51 49.60 17.18 13.92 13.92 7.81 52.14 26.83 21.31 61.19 16.99 58.95 36.03 661
@@
Method CIFAR10 CIFAR100 STL10 SUN397 Food101 Oxfordpet Flowers102 DTD EuroSAT Fgvc_Aircraft TinyImageNet ImageNet Caltech101 Caltech256 StanfordCars PCAM Average Time (s) CLIP 8.98 3.90 13.47 0.07 0.39 0.00 0.00 0.00 0.00 0.39 0.00 0.00 0.41 0.15 0.09 0.16 1.74 0 FT-TeCoA [32] 9.96 4.68 23.24 0.21 2.73 5.44 5.17 14.85 9.73 0.39 16.99 8.85 16.21 27.08 4.49 15.26 10.33 551 PMG-AFT (ours) 31.05 13.28 44.33 14.47 10.54 12.78 8.75 23.56 15.13 0.78 13.39 9.07 16.40 43.16 12.10 17.84 17.91 849
Implementation Details. We utilize the ViT-B/32 architecture of the CLIP model as the backbone and use the SGD optimizer to adversarially finetune the target model for 10 epochs. For fine-tuning, we update all parameters of the image encoder with a learning rate of 5e-5. For visual prompt, we introduce learnable parameters as prompts, adding them to both the image layer with the same dimensions as the original image and the token layer with a dimension of 100. The learning rate for the visual prompt is set at 40. We use $l_{\infty}$ norm PGD-2 [30] and PGD-10 [30] attacks with perturbation bounds of $\varepsilon=1/255,2/255$ and $4/255$ for both adversarial training and evaluation, respectively. For hyper-parameters, we set $\alpha$ to 1 and $\beta$ to 1. Each model in this paper is trained on two NVIDIA GeForce RTX 3090 GPUs.
@@ 4.3 Performance against AutoAttack
Figure 4: Zero-shot robust accuracy (a) and clean accuracy (b) under different perturbation bounds ($\varepsilon=1/255,2/255$ and $4/255$). We employ the same perturbation bound for training and testing.
@@ 4.6 Ablation Study
Method Avg Clean Acc (%) Avg Robust Acc (%) Output + KL 55.71 31.95 Output + $L_{2}$ 48.21 27.23 Feature + $L_{2}$ 38.32 22.90 Feature + COS 52.98 24.53